OpenStudy (anonymous):

System of Equations 3X-Y= -8 2X+ 5Y=-45 Simplify to get Y=? X=?

jigglypuff314 (jigglypuff314):

Hello and Welcome to OpenStudy! :) would you like to solve this using the elimination method or the substitution method?

OpenStudy (anonymous):

hey :) which ever method is easier, what do you think?

OpenStudy (anonymous):

how about substitution?

jigglypuff314 (jigglypuff314):

sure :) 3X-Y= -8 isolate the y -> y = 3x + 8 <- then plug in to the one below 2X+ 5Y=-45 so 2x + 5(3x + 8) = -45 distribute 2x + 15x + 40 = -45 solve for x

OpenStudy (anonymous):

x=5?

jigglypuff314 (jigglypuff314):

close :) there should be a negative I think

OpenStudy (anonymous):

ohh yeah i made a small mistake.. opps so x=-5

OpenStudy (anonymous):

whats next?

jigglypuff314 (jigglypuff314):

yep :) then plug it into one of the original equations so plug in x= -5 into y = 3x + 8 and solve for y

OpenStudy (anonymous):

y=-7?

jigglypuff314 (jigglypuff314):

correct! :)

jigglypuff314 (jigglypuff314):

you're welcome :)

jigglypuff314 (jigglypuff314):

you too :) glad I could help :)